# AeroSketch: Near-Optimal Time Matrix Sketch Framework for Persistent, Sliding Window, and Distributed Streams

### v1

@@ 3.1.5. General Unnormalized Model
1
2Function Initialize($d,N,R,\varepsilon$):
@@
3 $L\leftarrow\lceil\log_{2}R\rceil;M\leftarrow[]$
4 for $i=0$ to L1L-1 do
5    $M.\texttt{append}(\textsc{AeroSketch}.\textsc{Initialize}(d,\varepsilon,\theta=2^{i}\varepsilon N))$
6
7
8
9Function Update($\bm{a}_{i}$):
@@ 3 $L\leftarrow\lceil\log_{2}R\rceil;M\leftarrow[]$
10
11 for $j=0$ to L1L-1 do
12    while $len(M[j].\mathcal{S})>\frac{8}{\varepsilon}$ or $M[j].\mathcal{S}[0].t\leq i-N$ do
13       $M[j].\mathcal{S}$.Popleft()
14
15    if $\left\|\bm{a}_{i}\right\|_{2}^{2}\geq\theta$ then
16       $M[j].\mathcal{S}$.append($\bm{a}_{i}$, $s=M[j].\mathcal{S}[-1].t$, $t=i$)
17    else
18       $M[j].\textsc{Update}(\bm{a}_{i})$
19
20
21
22
23Function Query():
24 $j\leftarrow\min_{j}\{1\leq M[j].\mathcal{S}[0].s\leq T-N+1\}$
25 return $\mathsf{FD}$($M[j]$.Query(lb=TN,ub=Tlb=T-N,ub=T), $\bm{a}_{i}\in M[j].\mathcal{S}]$)
26
@@ 3.2. Analysis
3.2.1. Reducing the Failure Probability from Eq. 2 to Arbitrarily Small $\delta$
@@
We defer the detailed proof to the Appendix D.
[…]
2 for $h=1$ to $r$ do
3 $\bm{Z}_{h}^{\prime},\hat{\bm{\Sigma}}_{h}]\leftarrow\mathtt{simul\_iter}({\bm{C}^{\prime}_{i}}^{\top},k,\varepsilon_{\text{SI}}=0.2)$ // $\hat{\bm{\Sigma}}_{h}=\text{diag}\left(\hat{\sigma}_{h1},\hat{\sigma}_{h2}\dots,\hat{\sigma}_{hk}\right)$
4 $\xi_{h}\leftarrow\max_{\xi}\{\hat{\sigma}_{h\xi}^{2}\geq\theta\},\quad\bm{Z}_{h}\leftarrow\bm{Z}_{h}^{\prime}[:,1:\xi]$
5 $c_{h}\leftarrow 0$
6 for $g=1$ to $s$ do
7    $c_{hg}^{2}\leftarrow\texttt{power\_iteration}(\bm{C}_{i}^{\prime}-\bm{C}_{i}^{\prime}\bm{Z}_{h}\bm{Z}_{h}^{\top})$
8    $c_{h}\leftarrow\max(c_{h},c_{hg})$
9 if $c_{h}<c$ then $c\leftarrow c_{h},\quad\bm{Z}_{i}^{\prime}\leftarrow\bm{Z}_{h}^{\prime}$
10
@@ 8    $c_{h}\leftarrow\max(c_{h},c_{hg})$
$\begin{split}&\bm{C}_{i}^{\top}\bm{C}_{i}+\underline{\bm{Z}_{i}}\cdot\underline{\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}}+(\underline{\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}})^{\top}\underline{\bm{Z}_{i}}^{\top}-\underline{\bm{Z}_{i}}\cdot\underline{\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}}\cdot\underline{\bm{Z}_{i}}\cdot\underline{\bm{Z}_{i}}^{\top}\\ {{}={}}&(\bm{C}_{i}^{\prime}-\bm{C}_{i}^{\prime}\bm{Z}_{i}\bm{Z}_{i}^{\top})^{\top}(\bm{C}_{i}^{\prime}-\bm{C}_{i}^{\prime}\bm{Z}_{i}\bm{Z}_{i}^{\top})+\bm{Z}_{i}\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}+{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}\bm{Z}_{i}\bm{Z}_{i}^{\top}-\bm{Z}_{i}\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}\bm{Z}_{i}\bm{Z}_{i}^{\top}\\ {{}={}}&{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}\cancel{-\bm{Z}_{i}\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}}\cancel{-{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}\bm{Z}_{i}\bm{Z}_{i}^{\top}}+\cancel{\bm{Z}_{i}\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}\bm{Z}_{i}\bm{Z}_{i}^{\top}}\\ &\cancel{+\bm{Z}_{i}\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}}+\cancel{{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}\bm{Z}_{i}\bm{Z}_{i}^{\top}}-\cancel{\bm{Z}_{i}\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}\bm{Z}_{i}\bm{Z}_{i}^{\top}}={\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}.\end{split}$
@@
According to Theorem 2.2, the time cost of the Simultaneous Iteration corresponding to each $k_{i}$ is $O(d\ell k_{i}\log\ell)$. Therefore, the total time cost for all $i$ is:
@@ 4.1. Persistent Matrix Sketch (Problem 3)
1
2Function Initialize($d,\varepsilon$):
3 $\left\|\bm{A}_{0}\right\|_{F}^{2}\leftarrow 0$
4 AeroSketch.Initialize($d,\varepsilon,\theta=0$)
5
6Function Update($\bm{a}_{i}$):
7 $\left\|\bm{A}_{t}\right\|_{F}^{2}\leftarrow\left\|\bm{A}_{t-1}\right\|_{F}^{2}+\left\|\bm{a}_{i}\right\|_{2}^{2}$
@@
9 AeroSketch.Update($\bm{a}_{i}$)
10
11
12Function Query($t$):
13 return AeroSketch.Query(0, $t$)
14
@@
We defer the proof of Theorem 4.1 to Appendix E.
@@ 4.2. Tracking Approximate Matrix Sketch over Sliding Window (Problem 4)
1
2Function Initialize($d_{x},d_{y},\varepsilon,N,\theta$):
@@
3 $\ell\leftarrow\lceil\frac{2}{\varepsilon}\rceil,\quad\hat{\bm{A}}\leftarrow\bm{0}^{d_{x}\times l},\quad\hat{\bm{B}}\leftarrow\bm{0}^{d_{y}\times l}$
4 $\mathcal{S}\leftarrow$ an empty queue
5
6
7Function Update($\bm{x}_{i},\bm{y}_{i}$):
@@ 4 $\mathcal{S}\leftarrow$ an empty queue
8
9 while $\mathcal{S}[0].t+N\leq i$ do // snapshot expired
10    $\mathcal{S}.\textsc{popleft}()$
11
12 $\tilde{\bm{A}}\leftarrow\begin{bmatrix}\bm{A}_{i-1}^{\top}&\bm{a}_{i}\end{bmatrix}^{\top},\tilde{\bm{B}}\leftarrow\begin{bmatrix}\bm{B}_{i-1}^{\top}&\bm{b}_{i}\end{bmatrix}^{\top}$
13
14 if $\text{rows}(\tilde{\bm{A}})\geq 2\ell$ then // $\tilde{\bm{A}}$ has no zero rows
@@
16 else
17    $\bm{A}^{\prime}_{i},\bm{B}^{\prime}_{i}\leftarrow\tilde{\bm{A}},\tilde{\bm{B}}$
18
19 $\hat{\sigma}_{1}^{2},\hat{\bm{v}}_{1}\leftarrow\mathtt{power\_iteration}(\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top},k=O(\log_{2}\ell))$
20 if $\hat{\sigma}_{1}\geq\frac{\theta}{2}$ then // largest singular value
21    for $j=1$ to $\lceil\log_{2}\ell\rceil$ do
22       $k\leftarrow\min(2^{j},\ell)$
23       $\bm{Z}^{\prime},\hat{\bm{\Sigma}}]\leftarrow\mathtt{simul\_iter}(\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top},k,\varepsilon_{\text{SI}})$
24       $\bm{H}^{\prime},\hat{\bm{\Sigma}}]\leftarrow\mathtt{simul\_iter}({\bm{B}^{\prime}_{i}\bm{A}^{\prime}_{i}}^{\top},k,\varepsilon_{\text{SI}})$
25       if $\hat{\bm{\Sigma}}_{k}<\theta$ then
26          $k^{\prime}\leftarrow\arg\max_{k^{\prime}}\{\hat{\bm{\Sigma}}_{k^{\prime}}\geq\theta\}$
27          $\bm{Z}_{i}\leftarrow\bm{Z}^{\prime}{[:,1:k^{\prime}]},\quad\bm{H}_{i}\leftarrow\bm{H}^{\prime}{[:,1:k^{\prime}]}$
28          $\texttt{snapshot}\leftarrow(\bm{Z}_{i},\bm{Z}_{i}^{\top}\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top},\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top}\bm{H}_{i}^{\top},\bm{H}_{i})$
29          $\mathcal{S}.\textsc{append}\left(\texttt{snapshot},s=\mathcal{S}[-1].t+1,t=i\right)$
30          $\bm{A}_{i}\leftarrow\left(\bm{I}-\bm{Z}\bm{Z}^{\top}\right)\bm{A}_{i}^{\prime},\ \bm{B}_{i}\leftarrow\left(\bm{I}-\bm{H}\bm{H}^{\top}\right)\bm{B}_{i}^{\prime}$
31            break
32
33
34
35 else
36    $\bm{A}_{i}\leftarrow\bm{A}_{i}^{\prime},\ \bm{B}_{i}\leftarrow\bm{B}_{i}^{\prime}$
37
38
39Function Query():
40 $\bm{A}\bm{B}^{\top}\leftarrow\bm{A}_{T}\bm{B}_{T}^{\top}$
41 forall $(\bm{Z}_{i},\bm{Z}_{i}^{\top}\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top},\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top}\bm{H}_{i}^{\top},\bm{H}_{i})$ in $\mathcal{S}$ do
42    $\bm{A}\bm{B}^{\top}\leftarrow\bm{A}\bm{B}^{\top}+\bm{Z}_{i}\bm{Z}_{i}^{\top}\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top}+\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top}\bm{H}_{i}^{\top}\bm{H}_{i}-\bm{Z}_{i}\bm{Z}_{i}^{\top}\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top}\bm{H}_{i}^{\top}\bm{H}_{i}$
43 $\left[\bm{U},\bm{\Sigma},\bm{V}^{\top}\right]\leftarrow\text{SVD}(\bm{A}\bm{B}^{\top})$
44 ${\bm{A}}^{*}\leftarrow\bm{U}\sqrt{\max\left(\bm{\Sigma}-\sigma_{\ell}\bm{I},\bm{0}\right)},\quad{\bm{B}}^{*}\leftarrow\bm{V}\sqrt{\max\left(\bm{\Sigma}-\sigma_{\ell}\bm{I},\bm{0}\right)}$
45 return ${\bm{A}}^{*}[:,1:\ell],\quad{\bm{B}}^{*}[:,1:\ell]$
@@ 43 $\left[\bm{U},\bm{\Sigma},\bm{V}^{\top}\right]\leftarrow\text{SVD}(\bm{A}\bm{B}^{\top})$
We defer the proof of Theorem 4.2 to Appendix F.

### v2

@@ 3.1.5. General Unnormalized Model
Function Initialize($d,N,R,\varepsilon$):
@@
 $L\leftarrow\lceil\log_{2}R\rceil;M\leftarrow[]$
 for $i=0$ to L1L-1 do
    $M.\texttt{append}(\textsc{AeroSketch}.\textsc{Initialize}(d,\varepsilon,\theta=2^{i}\varepsilon N))$
   end for
Function Update($\bm{a}_{i}$):
@@
 for $j=0$ to L1L-1 do
    while $len(M[j].\mathcal{S})>\frac{8}{\varepsilon}$ or $M[j].\mathcal{S}[0].t\leq i-N$ do
       $M[j].\mathcal{S}$.Popleft()
      end while
    if $\left\|\bm{a}_{i}\right\|_{2}^{2}\geq\theta$ then
       $M[j].\mathcal{S}$.append($\bm{a}_{i}$, $s=M[j].\mathcal{S}[-1].t$, $t=i$)
      end if
    else
       $M[j].\textsc{Update}(\bm{a}_{i})$
      end if
   end for
Function Query():
 $j\leftarrow\min_{j}\{1\leq M[j].\mathcal{S}[0].s\leq T-N+1\}$
 return $\mathsf{FD}$($M[j]$.Query(lb=TN,ub=Tlb=T-N,ub=T), $\bm{a}_{i}\in M[j].\mathcal{S}]$)
@@ 3.2. Analysis
3.2.1. Reducing the Failure Probability from Eq. (2) to Arbitrarily Small $\delta$
@@
The detailed proof can be found in Appendix D.
$c\leftarrow+\infty,\quad r\leftarrow\log_{100}\frac{2}{\delta},\quad s\leftarrow 2\log_{3}\frac{2}{\delta}$
for $h=1$ to $r$ do
 $\bm{Z}_{h}^{\prime},\hat{\bm{\Sigma}}_{h}]\leftarrow\mathtt{simul\_iter}({\bm{C}^{\prime}_{i}}^{\top},k,\varepsilon_{\text{SI}}=0.2)$ // $\hat{\bm{\Sigma}}_{h}=\text{diag}\left(\hat{\sigma}_{h1},\hat{\sigma}_{h2}\dots,\hat{\sigma}_{hk}\right)$
 $\xi_{h}\leftarrow\max_{\xi}\{\hat{\sigma}_{h\xi}^{2}\geq\theta\},\quad\bm{Z}_{h}\leftarrow\bm{Z}_{h}^{\prime}[:,1:\xi]$
 $c_{h}\leftarrow 0$
 for $g=1$ to $s$ do
    $c_{hg}^{2}\leftarrow\texttt{power\_iteration}(\bm{C}_{i}^{\prime}-\bm{C}_{i}^{\prime}\bm{Z}_{h}\bm{Z}_{h}^{\top})$
    $c_{h}\leftarrow\max(c_{h},c_{hg})$
   end for
 if $c_{h}<c$ then $c\leftarrow c_{h},\quad\bm{Z}_{i}^{\prime}\leftarrow\bm{Z}_{h}^{\prime}$
end for
@@
$\begin{split}&\bm{C}_{i}^{\top}\bm{C}_{i}+\underline{\bm{Z}_{i}}\cdot\underline{\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}}+(\underline{\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}})^{\top}\underline{\bm{Z}_{i}}^{\top}-\underline{\bm{Z}_{i}}\cdot\underline{\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}}\cdot\underline{\bm{Z}_{i}}\cdot\underline{\bm{Z}_{i}}^{\top}\\ {{}={}}&(\bm{C}_{i}^{\prime}-\bm{C}_{i}^{\prime}\bm{Z}_{i}\bm{Z}_{i}^{\top})^{\top}(\bm{C}_{i}^{\prime}-\bm{C}_{i}^{\prime}\bm{Z}_{i}\bm{Z}_{i}^{\top})+\bm{Z}_{i}\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}+{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}\bm{Z}_{i}\bm{Z}_{i}^{\top}-\bm{Z}_{i}\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}\bm{Z}_{i}\bm{Z}_{i}^{\top}\\ {{}={}}&{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}\cancel{-\bm{Z}_{i}\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}}\cancel{-{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}\bm{Z}_{i}\bm{Z}_{i}^{\top}}\cancel{+\bm{Z}_{i}\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}\bm{Z}_{i}\bm{Z}_{i}^{\top}}\\ &\cancel{+\bm{Z}_{i}\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}}\cancel{+{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}\bm{Z}_{i}\bm{Z}_{i}^{\top}}\cancel{-\bm{Z}_{i}\bm{Z}_{i}^{\top}{\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}\bm{Z}_{i}\bm{Z}_{i}^{\top}}={\bm{C}_{i}^{\prime}}^{\top}\bm{C}_{i}^{\prime}.\end{split}$
@@
According to Theorem 2.2, the time cost of the Simultaneous Iteration corresponding to each $k_{i}$ is $O(d\ell k_{i}\log d)$. Therefore, the total time cost for all $i$ is:
@@ 4.1. Persistent Matrix Sketch (Problem 3)
Function Initialize($d,\varepsilon$):
 $\left\|\bm{A}_{0}\right\|_{F}^{2}\leftarrow 0$
 AeroSketch.Initialize($d,\varepsilon,\theta=0$)
Function Update($\bm{a}_{i}$):
 $\left\|\bm{A}_{t}\right\|_{F}^{2}\leftarrow\left\|\bm{A}_{t-1}\right\|_{F}^{2}+\left\|\bm{a}_{i}\right\|_{2}^{2}$
@@
 AeroSketch.Update($\bm{a}_{i}$)
Function Query($t$):
 return AeroSketch.Query(0, $t$)
@@
The detailed proof can be found in Appendix E.
@@ 4.2. Tracking Approximate Matrix Sketch over Sliding Window (Problem 4)
Function Initialize($d_{x},d_{y},\varepsilon,N,\theta$):
@@
 $\ell\leftarrow\lceil\frac{2}{\varepsilon}\rceil,\quad\hat{\bm{A}}\leftarrow\bm{0}^{d_{x}\times l},\quad\hat{\bm{B}}\leftarrow\bm{0}^{d_{y}\times l}$
 $\mathcal{S}\leftarrow$ an empty queue
Function Update($\bm{x}_{i},\bm{y}_{i}$):
@@
 while $\mathcal{S}[0].t+N\leq i$ do // snapshot expired
    $\mathcal{S}.\textsc{popleft}()$
   end while
 $\tilde{\bm{A}}\leftarrow\begin{bmatrix}\bm{A}_{i-1}^{\top}&\bm{a}_{i}\end{bmatrix}^{\top},\tilde{\bm{B}}\leftarrow\begin{bmatrix}\bm{B}_{i-1}^{\top}&\bm{b}_{i}\end{bmatrix}^{\top}$
 if $\text{rows}(\tilde{\bm{A}})\geq 2\ell$ then // $\tilde{\bm{A}}$ has no zero rows
@@
 else
    $\bm{A}^{\prime}_{i},\bm{B}^{\prime}_{i}\leftarrow\tilde{\bm{A}},\tilde{\bm{B}}$
   end if
 $\hat{\sigma}_{1}^{2},\hat{\bm{v}}_{1}\leftarrow\mathtt{power\_iteration}(\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top},k=O(\log_{2}\ell))$
 if $\hat{\sigma}_{1}\geq\frac{\theta}{2}$ then // largest singular value
    for $j=1$ to $\lceil\log_{2}\ell\rceil$ do
       $k\leftarrow\min(2^{j},\ell)$
       $\bm{Z}^{\prime},\hat{\bm{\Sigma}}]\leftarrow\mathtt{simul\_iter}(\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top},k,\varepsilon_{\text{SI}})$
       $\bm{H}^{\prime},\hat{\bm{\Sigma}}]\leftarrow\mathtt{simul\_iter}({\bm{B}^{\prime}_{i}\bm{A}^{\prime}_{i}}^{\top},k,\varepsilon_{\text{SI}})$
       if $\hat{\bm{\Sigma}}_{k}<\theta$ then
          $k^{\prime}\leftarrow\arg\max_{k^{\prime}}\{\hat{\bm{\Sigma}}_{k^{\prime}}\geq\theta\}$
          $\bm{Z}_{i}\leftarrow\bm{Z}^{\prime}{[:,1:k^{\prime}]},\quad\bm{H}_{i}\leftarrow\bm{H}^{\prime}{[:,1:k^{\prime}]}$
          $\texttt{snapshot}\leftarrow(\bm{Z}_{i},\bm{Z}_{i}^{\top}\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top},\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top}\bm{H}_{i}^{\top},\bm{H}_{i})$
          $\mathcal{S}.\textsc{append}\left(\texttt{snapshot},s=\mathcal{S}[-1].t+1,t=i\right)$
          $\bm{A}_{i}\leftarrow\left(\bm{I}-\bm{Z}\bm{Z}^{\top}\right)\bm{A}_{i}^{\prime},\ \bm{B}_{i}\leftarrow\left(\bm{I}-\bm{H}\bm{H}^{\top}\right)\bm{B}_{i}^{\prime}$
            break
         end if
      end for
 else
    $\bm{A}_{i}\leftarrow\bm{A}_{i}^{\prime},\ \bm{B}_{i}\leftarrow\bm{B}_{i}^{\prime}$
   end if
Function Query():
 $\bm{A}\bm{B}^{\top}\leftarrow\bm{A}_{T}\bm{B}_{T}^{\top}$
 forall $(\bm{Z}_{i},\bm{Z}_{i}^{\top}\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top},\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top}\bm{H}_{i}^{\top},\bm{H}_{i})$ in $\mathcal{S}$ do
    $\bm{A}\bm{B}^{\top}\leftarrow\bm{A}\bm{B}^{\top}+\bm{Z}_{i}\bm{Z}_{i}^{\top}\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top}+\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top}\bm{H}_{i}^{\top}\bm{H}_{i}-\bm{Z}_{i}\bm{Z}_{i}^{\top}\bm{A}^{\prime}_{i}{\bm{B}^{\prime}_{i}}^{\top}\bm{H}_{i}^{\top}\bm{H}_{i}$
   end forall
 $\left[\bm{U},\bm{\Sigma},\bm{V}^{\top}\right]\leftarrow\text{SVD}(\bm{A}\bm{B}^{\top})$
 ${\bm{A}}^{*}\leftarrow\bm{U}\sqrt{\max\left(\bm{\Sigma}-\sigma_{\ell}\bm{I},\bm{0}\right)},\quad{\bm{B}}^{*}\leftarrow\bm{V}\sqrt{\max\left(\bm{\Sigma}-\sigma_{\ell}\bm{I},\bm{0}\right)}$
 return ${\bm{A}}^{*}[:,1:\ell],\quad{\bm{B}}^{*}[:,1:\ell]$
@@
The detailed proof can be found in Appendix F.